A force which depends on x as

F(x) = 3x^2 - 4x moves a particle from

x1= 2.0 m to x2= 5.0 m. Calculate the

work done by the force.



1
Expert's answer
2021-01-29T16:56:40-0500

By the definition of the work done:


"F=\int F(x)dx.""W=\displaystyle\intop_{2.0}^{5.0} (3x^2-4x)dx=75 \ J."